# AED: Adaptable Error Detection for Few-shot Imitation Policy

### v2

@@ Abstract
We introduce a new task called Adaptable Error Detection (AED), which aims to identify behavior errors in few-shot imitation (FSI) policies based on visual observations in novel environments. The potential to cause serious damage to surrounding areas limits the application of FSI policies in real-world scenarios. Thus, a robust system is necessary to notify operators when FSI policies are inconsistent with the intent of demonstrations. This task introduces three challenges: (1) detecting behavior errors in novel environments, (2) identifying behavior errors that occur without revealing notable changes, and (3) lacking complete temporal information of the rollout due to the necessity of online detection. However, the existing benchmarks cannot support the development of AED because their tasks do not present all these challenges. To this end, we develop a cross-domain AED benchmark, consisting of 322 base and 153 novel environments. Additionally, we propose Pattern Observer (PrObe) to address these challenges. PrObe is equipped with a powerful pattern extractor and guided by novel learning objectives to parse discernible patterns in the policy feature representations of normal or error states. Through our comprehensive evaluation, PrObe demonstrates superior capability to detect errors arising from a wide range of FSI policies, consistently surpassing strong baselines. Moreover, we conduct detailed ablations and a pilot study on error correction to validate the effectiveness of the proposed architecture design and the practicality of the AED task, respectively.
@@ 1 Introduction
Our work makes three significant contributions: (1) We define a vital yet underexplored task called adaptable error detection (AED). (2) We introduce a novel benchmark and present PrObe, facilitating the research community to delve into this important area collectively. (3) We conduct thorough evaluations on our proposed benchmark, demonstrating the effectiveness of PrObe. It surpasses several baseline models and exhibits robustness across different policies. Our research is anticipated to serve as a key foundation for future real-world experiments in the field of FSI research.
2 Related Works
@@ 3 Preliminaries
As stated above, the expert and agent usually have different appearances or configurations. The DC policy $\pi(a\mid s,\mathcal{D})$ learns implicit mapping using current states $s$ and demonstrations $\mathcal{D}$ to compute agent actions $a$. Next, we present the unified architecture of DC policies and how they produce the action. When the observations and demonstrations are RGB-D images that only provide partial information, we assume that current history $h$ (observations from the beginning to the latest) and demonstrations are adopted as inputs.
@@ 4 Adaptable Error Detection (AED)
Let $c$ denote the category of agent’s behavior status, where $c\in\mathbf{C},\mathbf{C}=\{\textrm{normal},\textrm{error}\}$. When the agent with the trained FSI policy $\pi_{\theta}$ performs in a novel environment $e^{n}$, an adaptable error detector $\phi$ can access the agent’s rollout history $h$ and few expert demonstrations $\mathcal{D}^{n}$. It then predicts the categorical probability $\hat{y}$ of the behavior status for the latest state in the history $h$ by $\hat{y}=\phi(h,\mathcal{D}^{n})=P(c\mid enc(h,\mathcal{D}^{n})),$ where $enc$ denotes the feature encoder, and it may be $enc_{\phi}$ ($\phi$ contains its own encoder) or $enc_{\pi_{\theta}}$ (based on policy’s encoder). Next, let $y$ represent the ground truth probability, we evaluate $\phi$ via the expectation of detection accuracy over agent rollouts $X^{n}$ in all novel environments $E^{n}$:
@@
We explain the utilized protocol (shown in Figure 2) with a practical scenario: A company develops a home robot assistant. This robot can perform a set of everyday missions in customers’ houses (novel environments $E^{n}$), given a few demonstrations $\mathcal{D}^{n}$ from the customer. Before selling, the robot is trained in the base environments $E^{b}$ built by the company. In this scenario, both the agent’s and expert’s optimal policies $\pi^{\ast}$ are available during training, with an established assumption from the FSI society [8, 10, 11] that base environments are high controllable. Then, we can collect successful agent rollouts $X^{b}_{succ}$ and a few expert demonstrations $\mathcal{D}^{b}$ for all base environments  [^1] The successful agent rollouts are not expert demonstrations since thet might have different configurations.. Next, we also collect failed agent rollouts $X^{b}_{fail}$ by intervening in the agent’s $\pi^{\ast}$ at a critical timing (e.g., the moment to grasp objects) so that $X^{b}_{fail}$ can possess precise frame-level error labels.
With these resources, our utilized AED protocol consists of three phases: (1) The policy $\pi_{\theta}$ is optimized using successful agent rollouts $X^{b}_{succ}$ and few demonstrations $\mathcal{D}^{b}$. (2) The adaptable error detector $\phi$ is trained on agent rollouts $X^{b}_{succ}$, $X^{b}_{fail}$ and few demonstrations $\mathcal{D}^{b}$. Besides, the detector $\phi$ may use features extracted from policy $\pi_{\theta}$’s encoder, whose parameters are not updated in this phase. (3) The policy $\pi_{\theta}$ solves the task leveraging few demonstrations $\mathcal{D}^{n}$, and the detector $\phi$ monitors the policy’s behavior simultaneously. Notably, no agent rollouts are collected in this phase. Only a few demonstrations $\mathcal{D}^{n}$ are available since the agent is now in novel environments $E^{n}$.
@@ 5.1 Rollout Augmentation
To ensure a balanced number of successful and failed rollouts (i.e., $X^{b}_{succ}$ and $X^{b}_{fail}$), along with precise frame-level labels, we gather them using the agent’s optimal policy $\pi^{\ast}$ (with intervention). However, even if the policy $\pi_{\theta}$ is trained on $X^{b}_{succ}$, it inevitably diverges from agent’s optimal policy $\pi^{\ast}$ due to limited rollout diversity. Therefore, AED methods trained solely on the raw collected agent rollouts will be too sensitive to any subtle differences in the trained policy’s rollouts, leading to high false positive rates.
@@ 6.2 Analysis of Experimental Results
Due to the diverse nature of policy behaviors and the varying characteristics of baselines, achieving consistently superior results across all tasks is highly challenging. Nonetheless, according to Figure 4, our PrObe achieved the highest Top 1 counts, average ranking, and average performance difference in both metrics, indicating PrObe’s superiority among AED methods and its robustness across different policies. We attribute this to PrObe’s design, which effectively addresses the challenges in the AED task. The average performance difference is the average performance gain compared to the second best method across cases, calculated as $(\mathrm{score}-\mathrm{second\_best\_score})/\mathrm{second\_best\_score}$.
@@
In response to the second question, we summarize comprehensive ablations in the Appendix. First, Figure 11 indicates that the proposed rollout augmentation (RA) strategy increases the rollout diversity and benefits AED methods with temporal information. Second, Figure 12 demonstrates that PrObe’s design effectively improves performance. Third, Figure 13 illustrates PrObe’s performance stability by executing multiple experiment runs on a subset of tasks and computing the performance variance. Fourth, we examine how AED methods’ performance is influenced when receiving demonstrations with distinct qualities, as presented in Table 6. Lastly, we study the influence of viewpoint changes in Table 7. Regarding the third question, we provide the pilot study’s result on error correction in Table 8 to demonstrate the practicality of our AED task. Please refer to the corresponding paragraphs in the Appendix for exhaustive versions.
@@ 7 Conclusion
We point out the importance of monitoring policy behavior errors to accelerate the development of FSI research. To this end, we formulate the novel adaptable error detection (AED) task, whose three challenges make previous error detection methods infeasible. To address AED, we propose the novel Pattern Observer (PrObe) by detecting errors in the space of policy features. With the extracted discernible patterns and additional task-related knowledge, PrObe effectively alleviates AED’s challenges. It achieves the best performance, as confirmed by both our primary experiment and thorough ablation analyses. We also demonstrate PrObe’s robustness in the timing accuracy experiment and the learned embedding visualization. Finally, we provide the pilot study on error correction to reveal the practicality of our AED task. Our work is an essential cornerstone in developing future FSI research to conduct complex real-world experiments.
@@
This appendix presents more details of our work, following a similar flow to the main paper. First, Section A introduces the few-shot imitation (FSI) policies. Then, our PrObe’s details are provided in Section B. Besides, the strong AED baselines are depicted in Section C. Next, the designed multi-stage FSI tasks are exhaustively described in Section D, including task configurations, design motivations, and possible behavior errors. In addition, more experimental results are shown in Section E to support our claims. At last, we responsibly present our work’s limitations in Section F.
@@ D.1 General Task Attributes
  • Realistic: Our tasks support multi-source lighting, soft shadow, and complex object texture to make them closer to reality (cf. Figure 9). Existing benchmarks usually have no shadow (or only a hard shadow) and simple texture. Besides, all objects are are given reasonable weights. The aforementioned gravity and friction also make our missions more realistic.
@@ Appendix E Additional Experimental Results
This section reports more experimental results. We introduce the accuracy function used for determining an error prediction result in Section E.1. Furthermore, more details on the main experiment are provided in Section E.2, Next, all ablations are summarized in Section E.3. At last, the pilot study on error correction is shown in Section E.4.
@@ E.3 Ablation Study
Figure 12 illustrates the performance contribution of each component of PrObe, observed while monitoring the SCAN policy solving the Pick & Plcae task. The naive model (first column) removes the pattern extractor and utilizes a FC layer followed by an IN layer to transform history embeddings. Clearly, the designed components and objectives enhance performance, particularly with the addition of the pattern extractor into the model. We attribute this improvement to the fact that the extracted embedding patterns by our PrObe are more informative and discernible.
@@
E.4 Pilot Study on Error Correction
DCT policy w/ TaskEmbED w/ LSTMED w/ DCTED w/ PrObe
[…]
To further examine the practicality of our AED task, we conducted a pilot study on error correction. In this experiment, the FSI policy is paused after the AED methods detect an error. Then, a correction policy from [19], which resets the agent to a predefined safe pose, is applied. Finally, the FSI policy continues to complete the task. We conducted the study on the Press Button task, where errors are most likely to occur and be corrected. Besides, the correction policy is defined as moving to the top center of the table. We allowed the DCT policy to cooperate with the correction policy and four AED methods (SVDDED excluded), as summarized in Table 8.

### v4

@@ Abstract
We introduce a new task called Adaptable Error Detection (AED), which aims to identify behavior errors in few-shot imitation (FSI) policies based on visual observations in novel environments. The potential to cause serious damage to surrounding areas limits the application of FSI policies in real-world scenarios. Thus, a robust system is necessary to notify operators when FSI policies are inconsistent with the intent of demonstrations. This task introduces three challenges: (1) detecting behavior errors in novel environments, (2) identifying behavior errors that occur without revealing notable changes, and (3) lacking complete temporal information of the rollout due to the necessity of online detection. However, the existing benchmarks cannot support the development of AED because their tasks do not present all these challenges. To this end, we develop a cross-domain AED benchmark, consisting of 322 base and 153 novel environments. Additionally, we propose Pattern Observer (PrObe) to address these challenges. PrObe is equipped with a powerful pattern extractor and guided by novel learning objectives to parse discernible patterns in the policy feature representations of normal or error states. Through our comprehensive evaluation, PrObe demonstrates superior capability to detect errors arising from a wide range of FSI policies, consistently surpassing strong baselines. Moreover, we conduct detailed ablations and a pilot study on error correction to validate the effectiveness of the proposed architecture design and the practicality of the AED task, respectively. The AED project page can be found at https://aed-neurips.github.io/.
@@ 1 Introduction
Our work makes three significant contributions: (1) We define a vital yet underexplored task called Adaptable Error Detection (AED) and develop its associated benchmark to facilitate collective exploration by the research community. (2) We introduce PrObe, a novel method that monitors the policy’s behavior by retrieving patterns from its feature embeddings. (3) We conduct thorough evaluations on the proposed benchmark, demonstrating the effectiveness of PrObe. It surpasses several baselines and shows robustness across different policies. We anticipate that our research will serve as a key foundation for future real-world experiments in the field of FSI research.
2 Related Work
@@ 3 Preliminaries
As stated above, the expert and agent usually have different appearances or configurations. The DC policy $\pi(a\mid s,\mathcal{D})$ learns an implicit mapping using current states $s$ and demonstrations $\mathcal{D}$ to compute agent actions $a$. Next, we present the unified architecture of DC policies and how they produce the action. When the observations $o$ and demonstrations are RGB-D images that only provide partial information, we assume that the current history $h_{t}\coloneq(o_{1},o_{2},...,o_{t})$ and demonstrations $\mathcal{D}$ are adopted as inputs.
@@ 4 Adaptable Error Detection (AED)
Let $c$ denote the category of agent’s behavior status, where $c\in\mathbf{C},\mathbf{C}=\{\textrm{normal},\textrm{error}\}$. When the agent with the trained FSI policy $\pi_{\theta}$ performs in a novel environment $e^{n}$, an adaptable error detector $\phi$ can access the agent’s rollout history $h$ and a few expert demonstrations $\mathcal{D}^{n}$. It then predicts the categorical probability $\hat{y}$ of the behavior status for the latest state in the history $h$ by $\hat{y}=\phi(h,\mathcal{D}^{n})=P(c\mid enc(h,\mathcal{D}^{n})),$ where $enc$ denotes the feature encoder, and it may be $enc_{\phi}$ ($\phi$ contains its own encoder) or $enc_{\pi_{\theta}}$ (based on policy’s encoder). Next, let $y$ represent the ground truth probability, we evaluate $\phi$ via the expectation of detection accuracy over agent rollouts $X^{n}$ in all novel environments $E^{n}$:
@@
We explain the utilized protocol (shown in Figure 2) with a practical scenario: A company develops a home robot assistant. This robot can perform a set of everyday missions in customers’ houses (novel environments $E^{n}$), given a few demonstrations $\mathcal{D}^{n}$ from the customer. Before selling, the robot is trained in the base environments $E^{b}$ built by the company. In this scenario, both the agent’s and expert’s optimal policies $\pi^{\ast}$ are available during training, with an established assumption from the FSI society [8, 10, 11] that base environments are highly controllable. Then, we can collect successful agent rollouts $X^{b}_{succ}$ and a few expert demonstrations $\mathcal{D}^{b}$ for all base environments  [^1] The successful agent rollouts are not expert demonstrations since they might have different configurations.. Next, we also collect failed agent rollouts $X^{b}_{fail}$ by intervening in the agent’s $\pi^{\ast}$ at a critical timing (e.g., the moment to grasp objects) so that $X^{b}_{fail}$ can possess precise frame-level error labels.
With these resources, our utilized AED protocol consists of three phases: (1) The policy $\pi_{\theta}$ is optimized using successful agent rollouts $X^{b}_{succ}$ and a few demonstrations $\mathcal{D}^{b}$. (2) The adaptable error detector $\phi$ is trained on agent rollouts $X^{b}_{succ}$, $X^{b}_{fail}$ and a few demonstrations $\mathcal{D}^{b}$. Besides, the detector $\phi$ may use features extracted from policy $\pi_{\theta}$’s encoder, whose parameters are not updated in this phase. (3) The policy $\pi_{\theta}$ solves the task leveraging few demonstrations $\mathcal{D}^{n}$, and the detector $\phi$ monitors the policy’s behavior simultaneously. Notably, no agent rollouts are collected in this phase. Only a few demonstrations $\mathcal{D}^{n}$ are available since the agent is now in novel environments $E^{n}$.
@@ 5.1 Rollout Augmentation
To ensure a balanced number of successful and failed rollouts (i.e., $X^{b}_{succ}$ and $X^{b}_{fail}$), along with precise frame-level labels, we gather them using the agent’s optimal policy $\pi^{\ast}$ (with intervention). However, even if the policy $\pi_{\theta}$ is trained on $X^{b}_{succ}$, it inevitably diverges from the agent’s optimal policy $\pi^{\ast}$ due to limited rollout diversity. Therefore, AED methods trained solely on the raw collected agent rollouts will be too sensitive to any subtle differences in the trained policy’s rollouts, leading to high false positive rates.
@@ 6.2 Analysis of Experimental Results
Due to the diverse nature of policy behaviors and the varying characteristics of baselines, achieving consistently superior results across all tasks is highly challenging. Nonetheless, according to Figure 4, our PrObe achieved the highest Top 1 counts, average ranking, and average performance difference in both metrics, indicating PrObe’s superiority among AED methods and its robustness across different policies. We attribute this to PrObe’s design, which effectively addresses the challenges in the AED task. The average performance difference is the average performance gain compared to the worst method across cases, calculated as $(\mathrm{score}-\mathrm{worst\_score})/\mathrm{worst\_score}$.
@@
In response to the second question, we summarize comprehensive ablations in the Appendix. First, Figure 11 indicates that the proposed rollout augmentation (RA) strategy increases the rollout diversity and benefits AED methods with temporal information. Second, Figure 12 demonstrates that PrObe’s design effectively improves performance. Third, Figure 13 illustrates PrObe’s performance stability by executing multiple experiment runs on a subset of tasks and computing the performance variance. Fourth, we examine how AED methods’ performance is influenced when receiving demonstrations with distinct qualities, as presented in Table 7. Lastly, we study the influence of viewpoint changes in Table 8. Please refer to the corresponding paragraphs in the Appendix for exhaustive versions.
6.3 Pilot Study on Error Correction
To further examine the practicality of our AED task (the third question), we conducted a pilot study on error correction. In this experiment, the FSI policy is paused after the AED methods detect an error. Then, a correction policy from [19], which resets the agent to a predefined safe pose, is applied. Finally, the FSI policy continues to complete the task. We conducted the study on the Press Button task, where errors are most likely to occur and be corrected. Besides, the correction policy is defined as moving to the top center of the table. We allowed the DCT policy to cooperate with the correction policy and four AED methods (SVDDED excluded), as summarized in Table 1.
DCT policy w/ TaskEmbED w/ LSTMED w/ DCTED w/ PrObe
[…]
@@ 7 Conclusion
We point out the importance of monitoring policy behavior errors to accelerate the development of FSI research. To this end, we formulate the novel adaptable error detection (AED) task, whose three challenges make previous error detection methods infeasible. To address AED, we propose the novel Pattern Observer (PrObe) by detecting errors in the space of policy features. With the extracted discernible patterns and additional task-related knowledge, PrObe effectively alleviates AED’s challenges. It achieves the best performance, as confirmed by both our primary experiment and thorough ablation analyses. We also demonstrate PrObe’s robustness in the timing accuracy experiment and the learned embedding visualization. Additionally, we provide a pilot study on error correction, revealing the practicality of the AED task. Our work is an essential cornerstone in developing future FSI research to conduct complex real-world experiments.
Limitations and future work
We carefully discuss the limitations of our work in Section F of the Appendix, including online AED inference and real-world experiments. For future work, developing a unified AED method applicable to various tasks and policies is worth exploring. Additionally, discovering previously unseen erroneous behaviors remains an interesting and challenging avenue.
Acknowledgements
This work was supported in part by the National Science and Technology Council, Taiwan, under Grant NSTC 112-2634-F-002-006. We are grateful to MobileDrive and the National Center for High-Performance Computing. We also thank all reviewers and area chairs for their valuable comments and positive recognition of our work during the review process.
@@ Acknowledgements
This appendix presents more details of our work, following a similar flow to the main paper. First, Section A introduces the few-shot imitation (FSI) policies. Then, our PrObe’s details are provided in Section B. Besides, the strong AED baselines are depicted in Section C. Next, the designed multi-stage FSI tasks are exhaustively described in Section D, including task configurations, design motivations, and possible behavior errors. Lastly, more experimental results are shown in Section E to support our claims.
@@ D.1 General Task Attributes
  • Realistic: Our tasks support multi-source lighting, soft shadow, and complex object texture to make them closer to reality (cf. Figure 9). Existing benchmarks usually have no shadow (or only a hard shadow) and simple texture. Besides, all objects are given reasonable weights. The aforementioned gravity and friction also make our missions more realistic.
@@ Appendix E Additional Experimental Results
This section reports more experimental results. We introduce the accuracy function used for determining an error prediction result in Section E.1. Furthermore, more details on the main experiment are provided in Section E.2, Next, all ablations are summarized in Section E.3. At last, the pilot study on error correction is shown in Section 6.3.
@@ E.3 Ablation Study
Figure 12 illustrates the performance contribution of each component of PrObe, observed while monitoring the SCAN policy solving the Pick & Place task. The naive model (first column) removes the pattern extractor and utilizes a FC layer followed by an IN layer to transform history embeddings. Clearly, the designed components and objectives enhance performance, particularly with the addition of the pattern extractor into the model. We attribute this improvement to the fact that the extracted embedding patterns by our PrObe are more informative and discernible.